There were 430 people at a play. The admission price was $3 for adults and $1 for children. The admission receipts were $830. How many adults and how many children attended?
Answer by stanbon(75887)

There were 430 people at a play. The admission price was $3 for adults and $1 for children. The admission receipts were $830. How many adults and how many children attended?
----------------------
Equations:
Quantity: a + c = 430 tickets
Value::: 3a + c = 830 dollars
-------------------------------------
Subtract Quantity Eq from Value Eq and solve for "a":
2a = 400
a = 200 (# of adult tickets sold)
---
Since a+c = 430, c = 230 (# of child tickets sold)
======================
Cheers,
Stan H.